OpenStudy (anonymous):

Write the equation of the line that is parallel to the line y = -3x + 12 and passes through the point (-1, 6).

OpenStudy (anonymous):

parallel lines have the same slope follow equation y-y1=m(x-x1) y-6= -3(x-(-1) y-6=-3x-3 y=-3x+3 is the new equation that is parallel to y=-3x+12 Answer: y=-3x+3

OpenStudy (anonymous):

Thank you! Can you help me with one more?